Compute for the ideal length of a runway that is designed for airplanes that has a required speed of 35.0m/s for take-off at a rate of 3.5m/s2.
1
Expert's answer
2020-10-23T11:52:14-0400
"v^2=2as\\35^2=2(3.5)s\\s=175\ m"